The head of a fish is 1/3 its body. The tail of the fish is as long as its head and its body combined. The total length of the fish is 48cm. How long is each part of the fish?
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

The head of a fish is 1/3 its body. The tail of the fish is as long as its head and its body combined. The total length of the fish is 48cm. How long is each part of the fish?
-----------------
Head = 48/3 = 16 cm
Tail = 16 + 48 = 64 cm
------------
Who told you that?

Answer by stanbon(75887)

The head of a fish is 1/3 its body. The tail of the fish is as long as its head and its body combined. The total length of the fish is 48cm. How long is each part of the fish?
----
fish = head + body + tail
48 = head + body + tail
---
b = 3*h
t = h + 3h = 4h
----
Equation:
h + 3h + 4h = 48
8h = 48
head = h = 6 cm
body = 3h = 18 cm
tail = 4h= 24 cm
===================
Cheers,
Stan H.